# Selecting Critical Scenarios of DER Adoption in Distribution Grids Using Bayesian Optimization

### v1

@@ Results
Figure 5: Evolution of statistical stopping criteria $\tau_{n},\tau_{n}^{b}$ and $\tau_{n}^{l}$. The termination threshold is set to $\bar{\tau}=0.1$.
@@ IV-B Comparators
One baseline competitor is a brute force approach that sequentially evaluates every simulated scenario to identify the critical ones. In the above case study, this would require 27.6$\times$ more evaluations (9200/333) and would take 7.65$\times$ longer runtime after accounting for the overhead of our algorithm including the computation of the acquisition function and fitting GPs on all the bus and line objectives. Furthermore, the brute force approach lacks statistical analysis of criticality.
@@
Figure 6 illustrates how our methodology picks up bus-critical scenarios. The green curve counts the total number of critical scenarios in the search space $\mathcal{S}_{n}$, which recall increases in $n$. The orange curve shows the size of the subset of these that have been selected and evaluated by the BO, i.e. scenarios that are correctly identified as critical. Our algorithm is able to quickly identify critical scenarios shortly after they are added to the search space, cf. the fact that the orange curve closely tracks the green one. In contrast, the black staircase representing the number of critical scenarios found by the brute force approach lags significantly behind, unable to keep up with the growing size of $\mathcal{S}_{n}$, leading to a widening gap between the number of critical scenarios selected by our algorithm and those discovered through brute force.
@@
Figure 6: Number of bus-critical scenarios in the search space $\mathcal{S}_{n}$(green line) and how many of those scenarios are identified by Algorithm 1 and a greedy comparator as a function of time.
Figure 7: Simulated scenarios sorted by PV capacity aggregated over the networks p4rhs8 (top) and p3uhs72 (bottom) and colored by the maximum bus violation. Critical scenarios (evaluated by Algorithm 1) are indicated by a star.
@@
A second comparator approach is to rank simulated scenarios by total PV adoption due to the popular rule of thumb that more PV capacity causes more grid stress. In fact, in distribution planning, utilities consider the extreme cases of aggregated capacity (”high” and ”low” load or PV scenarios) as worst-cases to explore. This greedy approach evaluates the $N_{top}$ scenarios that have the highest adoption capacity. The top panel of Figure 7 shows all 9200 simulated scenarios sorted by their aggregated adopted PV, and colored by the maximum violation over the buses, with a star under scenarios that are critical with respect to bus objectives. Only 3 out of the top-25 scenarios are critical; in particular the scenario with the largest PV is not critical and only has line (but no bus) violations. With our method, we successfully find critical scenarios that the greedy approach misses. This result shows that the DER adoption scenarios critical for distribution grid planning are not necessarily those with extreme aggregated adoption patterns, usually explored by utilities, which highlights the importance of systematic locational search algorithms.
@@
Figure 8: Critical objective values found across 4 runs (green bars) of Algorithm 1 and a fixed database $\mathcal{S}_{H}$ with $H=25000$ scenarios. Orange bars show the maximum violation attained in the top $N_{top}=25$ scenarios by total PV. The horizontal lines show the maximum violation. Bus objectives $f_{k}$ (left $y$-axis) $k\in\{1,3,5,7,8\}$, and line objectives (right $y$-axis) $k\in\{18,19,30,31,50,56,66,151,193\}$.
@@ IV-C Second Feeder
To validate the performance of our algorithm, we consider a larger 439-bus network (feeder p3uhs27) with 9 bus groups and 300+ adopters (shown in left panel of Figure 9), see the right column of Table I. For this feeder, a stopping threshold of $\bar{\tau}=0.2$ is set and we use batch evaluations with $B=5$ scenarios. Results are similar to the first feeder p4rhs8. The search strongly targets scenarios leading to violations (over $95\%$ of evaluated scenarios result in bus voltage violations, and over $79\%$ in line flow ones). We recover 34 out of 36 bus-critical and 19 out of 20 line-critical scenarios. This shows the robustness of our methodology. The lower panel of Figure 7 shows the distribution of violations across the 21800 simulated scenarios, highlighting that violations are common (even for scenarios with lower total adopted PV capacity), but only a few scenarios are critical, and those are not necessarily the ones with the largest PV. Even if we evaluated the top 452 scenarios by total PV, we would miss a majority of the critical scenarios.
Figure 9: Left: Feeder p3uhs27 and bus groupings obtained using the Louvain community algorithm. Right: Spatial map of adopter relevance $\tilde{\theta}_{6,j}$ for representative bus objectives where each bubble represents an adopter $j$. Larger values (brighter colors) indicate stronger influence of the adopter on the respective objective. Circle size is proportional to the adopter’s PV system capacity.
@@ IV-D Algorithm Stability
To evaluate the stability of our search, we fix a large database $\mathcal{S}_{H}$ of $H=25000$ scenarios and run Algorithm 1 four times on feeder p4rhs8 where for each run $i=1,\ldots,4$ the search space $\mathcal{S}^{(i)}_{n}$ is constructed by randomly sampling without replacement from $\mathcal{S}_{H}$, with a different seed for every run. For each run, we record in Table II the number of function evaluations $n^{\star}$, the search space size at termination ($\mathcal{S}^{(i)}_{n}$), and the number of found bus- and line-critical scenarios. The maximum violation achieved by evaluated scenarios (i.e. the marginal Pareto front) for each of the 5+10 critical objectives is shown in Figure 8.
@@
Run Evaluated Simulated Critical (Buses) Critical (Lines)
1 392 10600 16 33
2 524 14950 15 31
3 398 11500 11 29
4 440 13000 15 29
TABLE II: Stability analysis of Algorithm 1. Each row corresponds to a different run where the search space is constructed by independent sampling from a fixed database of $H=25000$ scenarios.
@@ V-A Surrogate Performance
Figure 10: True stress objective ($x$-axis) vs. GP prediction, i.e. posterior mean ($y$-axis) for the critical bus objective $f_{8}(\mathbf{x})$, after evaluating $n=75$ (left) and $n=333$ scenarios (right) respectively. Vertical error bars indicate ±$\pm 2$ standard deviations around the mean. Orange points are in-sample scenarios and green points are out-of-sample. The rug plots indicate the distribution of the training outputs $\{f(\mathbf{x}),\mathbf{x}\in\mathcal{D}_{n}\}$.
It is instructive to illustrate the evolution of the GP surrogates $\widehat{f}_{k}(\cdot)$ over the BO iterations $n=n_{0},\ldots$. Because the search space $\mathcal{X}$ is high-dimensional, we cannot directly visualize $\hat{f}^{(n)}_{k}(\cdot)$; instead we show in Figure 10 a scatterplot of the predicted mean (and standard deviations) $\mu^{(n)}_{k}(\cdot)$ objective values against the true values $f_{k}(\cdot)$ at the initial step $n=n_{0}$ and at the last step $n=n^{\star}$ when the algorithm terminated. The predictions are shown for the already evaluated scenarios $\mathcal{D}_{n}$, as well as a (subset) of the current search set $\mathcal{S}_{n}$. The uncertainty of the predictions is indicated by the vertical bars that span $\mu^{(n)}_{k}(\mathbf{x})\pm 2\sigma^{(n)}_{k}(\mathbf{x})$. Note that for $\mathbf{x}\in\mathcal{D}_{n}$, there is almost no uncertainty due to the small $\sigma_{k}^{2}$.
@@
Accurate predictions would make $\mu^{(n)}_{k}(\mathbf{x})=f_{k}(\mathbf{x})$ lie on the diagonal. As $n$ increases, two effects occur. First, the algorithm focuses on scenarios that are close to critical, i.e. with $f_{k}(\mathbf{x})>0$. Therefore, it shifts learning to the “right”, cf. the rug plots in the left and right panels. Second, for $n=333$ the GPs have more training data and therefore yield more accurate (smaller prediction errors) and more precise (smaller $\sigma^{(n)}_{k}(\mathbf{x})$) fits. In conjunction, these two effects shift the training observations and asymmetrically improve accuracy, witness the right panel of Figure 10. The much improved quality of $\hat{f}_{k}(\cdot)$ in the top-right quadrant at $n=333$ compared to initial fit at $n=75$ shows that the surrogates are adept at learning individual $f_{k}$’s in their critical regions as the BO progresses.
V-B Most Relevant Adopters
@@
Our choice of the GP kernel furnishes a tool to interpret adopter relevance, i.e. to get information about which specific adopter buses are driving different bus and line stresses. Specifically, leveraging the separable nature of (5) the fitted relevance parameters $\boldsymbol{\theta}_{k}$ indicate how much changes in the adoption pattern $\mathbf{x}$ affect the GP predictions. High value for $\theta_{k,j}$ indicates that adoption by agent $j$ significantly impacts the stress in objective $k$. In that way, fitting the GP kernels as part of the BO provides the association between adopters and future grid stresses.
@@
Figure 11: Adopter relevance $\tilde{\theta}_{k,j}$ (brighter is higher) for the 12 bus objectives $f_{k}$, $k=1,\ldots,{12}$ (rows) in feeder p4rhs8 against adopters $j=1,\ldots,159$ (columns).
@@
We plot the relevance of each adopter $\tilde{\theta}_{k,j}:=1-e^{-\frac{\theta_{k,j}}{A}}$ in Figure 11 for feeder p4rhs8, with the bus objectives on the $y$-axis and the adopters on the $x$-axis, sorted by the partitions $\mathcal{B}_{p}$ to which they belong. The ARD feature of our GPs endogenously identifies the relevant adopters for each objective, shown by the brighter colors. The sparsity of the heatmap implies that for each bus objective, only a few adopters are relevant and there is a low-dimensional subspace of $\mathcal{X}$ with “active” dimensions. This sparsity helps the GPs to focus on these dimensions to model the stress values. The right panel of Figure 9 visualizes $\tilde{\theta}_{7,j}$ for objective $f_{7}$ (based on buses indicated by the orange points) for feeder p3uhs27 spatially overlaid on the grid network. We find that the resulting most-relevant “active” adopters are consistent with grid topology, corresponding to buses that are in the cluster $\mathcal{B}_{7}$ defining $f_{7}$, or to nearby adopters with large PV capacities.

### v2

@@ Results
Figure 5: Evolution of statistical stopping criteria $\tau_{n}^{b}$ for buses and $\tau_{n}^{l}$ for lines as a function of number of evaluations $n$. The overall stopping criterion $\tau_{n}=\max(\tau^{b}_{n},\tau^{l}_{n})$ is triggered at the threshold $\bar{\tau}=0.1$.
@@ IV-B Comparators
[…]
@@
Figure 6 illustrates how our methodology picks up bus-critical scenarios. The green curve counts the total number of critical scenarios in the search space $\mathcal{S}_{n}$, which increases in $n$. The orange curve shows the size of the subset of these that have been selected and evaluated by the BO, i.e. scenarios that are correctly identified as critical. Our algorithm is able to quickly identify critical scenarios shortly after they are added to the search space, cf. the fact that the orange curve closely tracks the green one. In contrast, the black staircase representing the number of critical scenarios found by the brute force approach lags significantly behind, unable to keep up with the growing size of $\mathcal{S}_{n}$, leading to a widening gap between the number of critical scenarios selected by our algorithm and those discovered through brute force.
@@
Figure 6: Number of bus-critical scenarios in the search space $\mathcal{S}_{n}$(green line) and how many of those scenarios are identified by Algorithm 1 (orange line) and a greedy comparator (in black) as a function of time. The number of critical scenarios found by a naïve comparator that evaluates the $N_{top}=25$ scenarios by total PV is the magenta curve. When Algorithm 1 stops after 7.5 minutes, the greedy approach has only identified 4 critical scenarios.
Figure 7: Bus violations within the simulated PV adoption scenarios for feeders p4rhs8 (top, 9200 scenarios) and p3uhs27 (bottom, 21800 scenarios), sorted left-to-right by their total installed PV capacity. Each vertical bar corresponds to a scenario, with color indicating the maximum bus voltage violation observed under that scenario. Adoption scenarios identified as critical by Algorithm 1 (i.e., belonging to the Pareto set of violations) are marked with a star.
@@
A second comparator approach is to rank simulated scenarios by total PV adoption and compare the critical scenarios obtained by the proposed search algorithm with the existing utility practice, which consists of selecting the extreme cases of aggregated capacity (“high” and “low” load or PV scenarios) as worst-cases to explore. This naïve approach evaluates the $N_{top}$ scenarios that have the highest adoption capacity. Figure 6 shows that this naïve approach (with $N_{top}=25$) misses the majority of critical scenarios that our algorithm finds and barely improves upon greedily evaluating every simulated scenario. The top panel of Figure 7 further shows the distribution of the 9200 simulated scenarios sorted by their aggregated adopted PV, and colored by the maximum violation over the buses. Notably, most scenarios with the largest PV adoption are not critical, while other scenarios with relatively low PV capacity are critical (indicated with a star) due to specific locational adoption patterns. For example, high PV penetration throughout a feeder can lead to overvoltages at nodes distant from the substation while alleviating overloading in lines closer to it. However, a moderate PV penetration at the end of the feeder can still cause the same overvoltages but may not be sufficient to alleviate overloading, resulting in a more challenging planning scenario. This highlights the importance of systematic locational search algorithms and explains how, with our method, we successfully find critical scenarios that the greedy and naive approaches miss.
@@
Figure 8: Critical objective values found across 4 runs (green bars) of Algorithm 1 and a fixed database $\mathcal{S}_{H}$ with $H=25000$ scenarios. Orange bars show the maximum violation attained in the top $N_{top}=25$ scenarios by total PV. The horizontal lines show the maximum violation in this database. The critical bus objectives are $f_{k}$ (left $y$-axis) $k\in\{1,3,5,7,8\}$, and the critical line objectives (right $y$-axis) are $k\in\{18,19,30,31,50,56,66,151,193\}$.
@@ IV-C Second Feeder
To validate the performance of our algorithm, we consider a larger 439-bus network (feeder p3uhs27) with 9 bus groups and 300+ adopters (shown in left panel of Figure 9), see the right column of Table I. For this feeder, a stopping threshold of $\bar{\tau}=0.2$ is set and we use batch evaluations with $B=5$ scenarios. Results are similar to the first feeder p4rhs8. The search strongly targets scenarios leading to violations (over $95\%$ of evaluated scenarios result in bus voltage violations, and over $79\%$ in line flow ones). We recover 34 out of 36 bus-critical and 19 out of 20 line-critical scenarios. This shows the robustness of our methodology. The lower panel of Figure 7 shows the distribution of the total adopted PV of the 21800 simulated scenarios and color-coded by the maximum bus violation. The Figure highlights that violations are common (even for scenarios with lower total adopted PV capacity), but only a few scenarios are critical, and those are not necessarily the ones with the largest PV. Even if we evaluated the top 452 scenarios by total PV, we would miss a majority of the critical scenarios.
Figure 9: Left: Feeder p3uhs27 and bus groupings obtained using the Louvain community algorithm. Right: Spatial map of adopter relevance $\tilde{\theta}_{1,j}$ for a representative bus objective $f_{1}$. Adopters $j$ marked with an orange dot are buses that are part of objective $f_{1}$. Larger values (brighter colors) indicate stronger influence of the adopter on the respective objective. Circle size is proportional to the adopter’s PV system capacity.
@@ IV-D Algorithm Stability
[…]
@@
Simulated Evaluated Critical Critical
Run Bus scens Line scens
1 10600 392 16 33
2 14950 524 15 31
3 11500 398 11 29
4 13000 440 15 29
TABLE II: Stability analysis of Algorithm 1. Each row is a different run with the search space constructed by independent sampling from a fixed database of $H=25000$ scenarios.
Next, we perform a sensitivity analysis for four key parameters in Algorithm 1: the number of candidate scenarios $M$ at each BO step, the number of samples $N$ to compute $\alpha_{ND}$ in (8), the stopping threshold $\bar{\tau}$, and the number of new scenarios simulated at each BO step $N_{\text{expand}}$. To control for the intrinsic randomness due to scenario simulation, all the reported values are averages over 10 runs of Algorithm 1; we moreover list the respective standard deviations. Table III reports the total number of simulated scenarios, the number of evaluations (equivalent to computational time), and the number of found critical scenarios for the bus and line objectives. we report the average number of simulated, evaluated, bus-critical and line-critical scenarios. The discussion below is relative to the baseline parameters from Table I.
While the number of scenarios simulated by our algorithm is by design sensitive to above tuning parameters, robustness is reflected in the stable number of evaluated scenarios, as well as a consistent number of found critical scenarios. Performance is little impacted by $N$, although setting it too low leads to a larger variance in all key metrics, i.e., a less stable performance, as the acquisition function becomes too noisy. The number of candidates in the stopping criterion, $M$, is also benign, though it must be sufficiently large to prevent the algorithm from terminating prematurely due to a lack of promising scenarios considered when computing the stopping criterion. The algorithm is most sensitive to $N_{e}=N_{\text{expand}}$ that controls the growth of the search space. Setting $N_{\text{expand}}$ too low causes the algorithm to exhaust the search space and stop prematurely. Taking it too large, causes the algorithm to be overwhelmed by the number of potential candidates with the result that the stopping criterion decreases very slowly. Finally, the threshold $\bar{\tau}$ directly controls the stopping rule, and a large $\tau$ causes early termination, missing some critical scenarios.
In terms of violation magnitudes, we ran a statistical test that confirmed that the distribution of maximum found violations is statistically the same between the baseline and the variants in Table III, so the algorithm consistently finds the worst violations of each objective. Overall, the stable number of evaluations ($|\mathbf{X}_{n^{\star}}|\in[300,370]$) and of the critical scenarios (141614-16 for buses 222422-24 for lines) discovered, indicates the robustness of our approach; the exception being the rows with $N=10$ (very large variance across runs) and $M=500$, $N_{\text{expand}}=50$ and $\bar{\tau}=0.5$ (insufficient number of evaluations leading to some critical scenarios missed).
[…]
TABLE III: Sensitivity analysis of Algorithm 1. Each row corresponds to the average of 10 runs on a fixed database $\mathcal{S}_{H}$ of $H=25000$ scenarios, with $\pm$ indicating the respective standard deviation across runs. Baseline parameters corresponding to the results in Table I are in bold.
Another algorithm parameter that ought to be mentioned is the number of bus groups $P$ that determines the number of bus objectives. This parameter represents the number of voltage control areas in the network, which are often a small number and very specific of each feeder. Similarly, the violation threshold $\bar{s}$ (Line 5 of Algorithm 1) is another user-defined parameter that is network-dependent. One way to pick $\bar{s}$ is to look at the distribution of objective values on the set of initial evaluated scenarios.
@@ V-A Surrogate Performance
Figure 10: True stress objective ($x$-axis) vs. GP prediction, i.e. posterior mean ($y$-axis) for the critical bus objective $f_{8}(\mathbf{x})$, after evaluating $n=75$ (left) and $n=333$ scenarios (right) respectively. Vertical error bars indicate $\pm 2$ standard deviations around the mean. Orange points are in-sample scenarios and green points are out-of-sample. The rug plots indicate the distribution of the training outputs $\{f(\mathbf{x}),\mathbf{x}\in\mathcal{D}_{n}\}$.
[…]
@@
Accurate predictions would make $\mu^{(n)}_{k}(\mathbf{x})=f_{k}(\mathbf{x})$ lie on the diagonal. As $n$ increases, two effects occur. First, the algorithm focuses on scenarios that are close to critical, i.e. with $f_{k}(\mathbf{x})>0$. Therefore, it shifts learning to the “right”, cf. the rug plots in the left and right panels. Second, for $n=333$ the GPs have more training data and therefore yield more accurate (smaller prediction errors) and more precise (smaller $\sigma^{(n)}_{k}(\mathbf{x})$) fits. In conjunction, these two effects shift the training observations and asymmetrically improve accuracy, as witnessed in the right panel of Figure 10. The much improved quality of $\hat{f}_{k}(\cdot)$ in the top-right quadrant at $n=333$ compared to initial fit at $n=75$ shows that the surrogates are adept at learning individual $f_{k}$’s in their critical regions as the BO progresses.
V-B Kernel Selection
The performance of the GP surrogate in Algorithm 1 hinges on the choice of its kernel. Standard kernels, such as the Matèrn or radial basis function (RBF), assume continuous inputs and are ill-suited for binary vectors modeling our adoption scenarios. The Hamming distance we employ is the natural choice to measure distance between scenarios. An alternative is to employ embedding techniques to project the high-dimensional discrete $\mathbf{x}$ into a lower-dimensional continuous space, where an RBF or Matèrn kernel can then be applied. Embedding can be done via random projection, for example the REMBO algorithm [47]; via dictionary embedding based on Hamming distance to an exogenously specified basis set, or via variational autoencoders.
Embedding is conceptually promising for our problem since only a subset of adopters is expected to influence each bus or line objective. Thus, the effective dimensionality of our problem is likely much lower than the input dimension. Motivated by this, we compare our categorical kernel to two prominent embedding approaches. First, we implemented REMBO [47] and vary the embedding dimension and the resampling frequency of the random projection matrix, and test both RBF and Matèrn kernels. Second, we applied the dictionary embedding method (BODi) [48], evaluating both the binary-wavelet dictionary and a variant of our own, where at each step the scenarios in the Pareto set are used as basis vectors. For compactness, we report only the best-performing runs. Importantly, we emphasize that our use of these kernels is limited to surrogate modeling; we still evaluate our acquisition function on a subset of the search space.
Figure 11: Number of bus-critical scenarios in the search space $\mathcal{S}_{n}$ (black curve) and how many of those scenarios are identified by Algorithm 1 using four different GP kernels, cf. Table IV. Each run is on the same search space but has its own stopping criterion, leading to different termination times.
Figure 11 reports the number of critical scenarios identified when running Algorithm 1 with different kernels. REMBO was manually terminated after 500 evaluations because its stopping criterion remained far from $\bar{\tau}$ and showed no indication of halting. BODi failed to recognize any scenarios as potentially critical and stopped prematurely after fewer than 50 evaluations, despite requiring a larger number of initial training points. For both REMBO and BODi, we fixed the embedding dimension at 20 (compared to $A=159$ for the categorical kernel), which substantially reduced GP fitting costs and which explains the shorter runtime for REMBO, despite having more function evaluations. The poor performance of REMBO and BODi stems from their apparent inability to capture the dependence of bus and line voltages on adoption scenarios. Table IV summarizes out-of-sample prediction errors across 500 scenarios, comparing our categorical kernel (5) the embedding-based alternatives. For every bus objective, the $\kappa_{k}$ in (5) achieves consistently lower RMSE and outperforms both REMBO and BODi. Similar results hold for line objectives, underscoring the advantage of directly modeling the binary structure of the input space.
REMBO BODi (wavelet) BODi (Pareto set) Cat. (Eqn (5))
$f_{1}$ 1.62 1.12 1.17 0.43
$f_{3}$ 1.63 0.99 0.95 0.41
$f_{5}$ 0.94 0.6 0.58 0.36
$f_{7}$ 0.8 0.69 0.65 0.15
$f_{8}$ 1.57 1.01 1.12 0.33
TABLE IV: Predictive accuracy of the five critical bus objectives for different choices of GP kernels. We report predictive RMSE $(\times 10^{-2})$ over 500 unevaluated scenarios randomly selected from the search space after evaluating $n=333$ scenarios selected by Algorithm 1.
V-C Most Relevant Adopters
@@
Our choice of the categorical GP kernel furnishes a tool to interpret adopter relevance, i.e. to get information about which specific adopter buses are driving different bus and line stresses. Specifically, leveraging the separable nature of (5) the fitted relevance parameters $\boldsymbol{\theta}_{k}$ indicate how much changes in the adoption pattern $\mathbf{x}$ affect the GP predictions. High value for $\theta_{k,j}$ indicates that adoption by agent $j$ significantly impacts the stress in objective $k$. In that way, fitting the GP kernels as part of the BO provides the association between adopters and future grid stresses.
@@
Figure 12: Heatmap of adopter relevance $\tilde{\theta}_{k,j}$ (brighter is higher) in feeder p4rhs8. Rows represent bus objectives $f_{k}$, $k=1,\ldots,{12}$, columns are the adopters $j=1,\ldots,159$. Vertical white lines separate adopters according to the bus groups they belong to.
@@
We plot the relevance of each adopter $\tilde{\theta}_{k,j}:=1-e^{-\frac{\theta_{k,j}}{A}}$ in Figure 12 for feeder p4rhs8, with the bus objectives on the $y$-axis and the adopters on the $x$-axis, sorted by the partitions $\mathcal{B}_{p}$ to which they belong. The ARD feature of our GPs endogenously identifies the relevant adopters for each objective, shown by the brighter colors. The sparsity of the heatmap implies that for each bus objective, only a few adopters are relevant and there is a low-dimensional subspace of $\mathcal{X}$ with “active” dimensions. This sparsity helps the GPs to focus on these dimensions to model the stress values. The right panel of Figure 9 visualizes $\tilde{\theta}_{1,j}$ for objective $f_{1}$ (based on buses indicated by the orange points) for feeder p3uhs27 spatially overlaid on the grid network. We find that the resulting most-relevant “active” adopters are consistent with grid topology, corresponding to buses that are in the cluster $\mathcal{B}_{1}$ defining $f_{1}$, or to nearby adopters with large PV capacities.
The identification of these relevant adopters provides utilities with a group of consumers that are likely to trigger grid investments through their adoption decisions. This information can be used to schedule grid upgrades in response to PV interconnection requests from this group. It can also offer useful insights on where to incentivize load growth or procure local flexibility services to defer such upgrades.